Opened 12 years ago

Last modified 11 years ago

#356 closed flight processing

EUFAR10/03, flight day 196/2010,Lisbon — at Version 23

Reported by: ella Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by benj)

Data location: ~arsf/arsf_data/2010/flight_data/western_europe/EUFAR10_03-2010_196_Lisbon

Data arrived from ARSF via network transfer on 2nd/08/2010.

Scientific objectives:
Objective 1.- Development of a reliable tool to estimate coastal water quality indicators by using remote sensing
techniques, both airborne and satellite.
1.1. | Development of a set of semi-empirical algorithms based on hyper-spectral airborne and in-situ data for
estimating the concentrations of the water constituents;
1.2. | Migration of the developed algorithms into the satellite workflow.

Priority: a4h

PI: Jose da Silva


  • Eagle
  • Hawk
  • Leica LIDAR
  • RCD Camera (16/09/10)

Change History (24)

comment:1 Changed 12 years ago by ella

  • Description modified (diff)
  • Priority changed from alpha 4 medium to alpha 4 high

comment:2 Changed 12 years ago by ella

Starting navigation processing.

comment:3 Changed 12 years ago by jb362

IPAS navigation processing done.

comment:4 Changed 12 years ago by ella

Waiting for logsheet.

Applanix processing was problematic because the flight was 100km away from the nearest base station.
POSproc failed to create an Sbet file.

comment:5 Changed 12 years ago by jb362

Starting camera processing.

comment:6 Changed 12 years ago by emca

Starting LiDAR processing

Changed 12 years ago by emca

line 06

comment:7 Changed 12 years ago by emca

There are six LiDAR lines for this project.

Lines 1 and 2 are very short, contain very little data (approx 30% first returns) and are very noisy. From the leica flightline log it it looks like there was also an issue with the range correction during the capture of these lines.

Lines 3 and 4 contain no useful data.

Line 5 is partially over water and contains a 90 degree turn.

Line 6 consists of four distinct lines, including the turns in between (see attached screenshot). This may need to be separated into 4 lines.

There is little overlap between lines (just the cross of Line 6 with itself) which will make correcting for roll difficult.

Going to wait until next week when ARSF Ops return and we can get the logsheet before proceeding any further with this project.

comment:8 Changed 12 years ago by jb362

Started camera processing for this project but in the kml file it's very hard to see which images will be needed and which won't. This is because it the camera wasn't turned off between flight lines. This processing will therefore be held until we have the logsheet for this project. If I have left before we do get the logsheet iopa will be picking up the camera processing.

comment:9 Changed 12 years ago by jb362

Also, all the .tif files for the camera processing have already been created.

comment:10 Changed 12 years ago by emca

Call from Phil to say there are still looking for the logs sheets for this project as well as for Lisbon flight day 197

Both projects were flown on an ad hoc basis. They were unable to turn the LiDAR off during the flight as they were flying over water but the hyperspectral was switched off between lines to conserve disk space. As a result there will not be a match between the eagle/hawk and the LiDAR lines.

comment:11 Changed 12 years ago by emca

Logsheet has arrived and been saved in ~arsf/arsf_data/2010/flight_data/archive-overflow/EUFAR10_03-2010_196_Lisbon/admin

comment:12 Changed 12 years ago by jb362

Iopa will be taking over this camera project.

comment:13 Changed 12 years ago by ella

Starting camera delivery check.

comment:14 Changed 12 years ago by ella

Camera check complete.
Two photos have been removed and the read_me and kml file have been updated.
Ready to deliver.

comment:15 Changed 12 years ago by iopa

  • Description modified (diff)

RCD camera was dispatched today, 16th September 2010.

comment:16 Changed 12 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Taking over hyper-spectral processing for this project.

comment:17 Changed 12 years ago by knpa

Issue with applanix - raw data in two parts with a gap in the middle. This gap covers one of the lines (Gen3).

Also, cannot find eagle data for first two lines (Gen 1/2) and (Gen 3).

comment:18 Changed 12 years ago by knpa

  • Description modified (diff)

Move to repository, can be rsynced.

comment:19 Changed 12 years ago by knpa

Phone call from Phil to Ben:
The data collection for this project starts at 11:39 and earlier lines are tests.

That excludes Gen1/2 and Gen3, meaning that this project now has 5 lines containing eagle and hawk data that will be processed.

comment:20 Changed 12 years ago by knpa

SCTs impossible to correct on this project as almost all lines are over water.
H6 had no sync but managed to correct gross sync issues by matching water patterns with eagle.


Eagle Hawk
Line 1 0 0
Line 2 0 0
Line 3 0 0
Line 4 0 0
Line 5 0 1.5 (sync)

comment:21 Changed 12 years ago by knpa

Correction: above Hawk 5 value should be 0.15 NOT 1.5.

comment:22 Changed 12 years ago by knpa

First 12 bands of Hawk lev1s are blank. The raws are fine. Appears to be an issue with the bandset labels in the header files, they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:23 Changed 12 years ago by benj

  • Description modified (diff)

Corrected PI (recorded wrong on data unpacking)

Note: See TracTickets for help on using tickets.